I wish to fortify my 16% wine by adding 90% Everclear alcohol. I want 2 liters of 40% finished product. how much of each should i use

Answer by Alan3354(69443)

I wish to fortify my 16% wine by adding 90% Everclear alcohol. I want 2 liters of 40% finished product. how much of each should i use
------------------
w + e = 2000 cc (total liquid)
16w + 90e = 40*2000 (total alcohol)
e = 2000 - w
----
16w + 90*(2000-w) = 80000
16w + 180000 - 90w = 80000
-74w = -100000
w = 1351.35 cc of wine
e = 648.65 cc of Everclear
==============